in 1975 the price of a frosty refrigerator was $480. a codwell refrigerator sold for $545. the price of a frosty went up $38 annually while the codwell went up $25 annually. in what year did the two refrigerators sell for the same price?
Answer by htmentor(1343)

We can write the cost of the frosty fridge as:
F=480+38t
And the cost of the codwell fridge as:
C=545+25t
We want to find t when F=C:
480+38t=545+25t -> 13t=65 -> t = 5
Ans: t = 5 years